# A spectral method for multi-view subspace learning using the product of projections

### v1

@@ 5.2. Nutrigenomic mice data
Figure 3. Left: Scree plot for gene expression and lipid content data. Blue point denotes the cutoff suggested by Gavish & Donoho (2014). Red point indicates our suggested cutoff. Middle / Right: Spectrum of the product of projection matrices onto each view for two different marginal ranks. The blue region indicates the random direction bound, $[0,\lambda_{+}^{1/2}]$, used to filter out the noise directions. The green region is given by $[1-\widehat{\varepsilon}_{1},1]$ where $\widehat{\varepsilon}_{1}$ is a bootstrap estimate of $\varepsilon_{1}$.
Table 3. Mean genotype and diet misclassification rates in percentages together with the minimum principle angle between estimated individual subspaces. We denote principle components analysis by PCA, Lock et al. (2013) as JIV, Feng et al. (2018) as AJI, Gaynanova & Li (2019) as SLI, Shu et al. (2019) as DCC, Park & Lock (2020) as UNI, and our method as PPD.
@@ Appendix A Techinical details
A.1. Proofs of main results
@@
[…]
@@
$\lambda_{n}(Y^{\top}Y)+\lambda_{r}(X^{\top}X)$ $\leq\lambda_{r}(X^{\top}X+Y^{\top}Y)\leq\lambda_{1}(Y^{\top}Y)+\lambda_{r}(X^{\top}X)$
@@
$\lambda_{r}(X^{\top}X+Y^{\top}Y)\leq\min\left\{\lambda_{r}(X^{\top}X)+\lambda_{1}(Y^{\top}Y),\lambda_{1}(X^{\top}X)+\lambda_{r}(Y^{\top}Y)\right\}$
Recall that $\sigma^{2}_{r}(A)=\lambda_{r}(A^{\top}A)$. Since $\lambda_{n}(Y^{\top}Y)\geq 0$, $\lambda_{n}(X^{\top}X)\geq 0$, we get (1). Similarly, since $\lambda_{1}(Y^{\top}Y)\geq 0$, $\lambda_{1}(X^{\top}X)\geq 0$, taking $r=1$ yields (2). ∎
Theorem A.2 (Spectrum of the product of projection matrices).
Let $\widehat{M}:=P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$. Suppose that $\operatorname*{rank}(X_{1}),\operatorname*{rank}(X_{2})>0$. Define
$\epsilon_{1}:=\left\|P_{{\operatorname*{col}(X_{1})}}\left(\Delta_{1}+\Delta_{2}+\Delta_{1}\Delta_{2}\right)P_{{\operatorname*{col}(X_{2})}}\right\|_{2},$ $\epsilon_{2}:=\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{2}+\Delta_{1}P_{{\operatorname*{col}(X_{2})}}+\Delta_{1}\Delta_{2}\right\|_{2}$
with $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$. Then the spectrum of $\widehat{M}$ consists of:
  1. (1)
    a group of $\dim(\mathcal{J})$ singular values in $\left[1-\epsilon_{1}\vee 0,1\right];$
  2. (2)
    a group of $\mathrm{rank}\left(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right)$ singular values in $\left[\left(\sigma_{\min}\left(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right)-\epsilon_{1}\right)\vee 0,\left(\sigma_{\max}\left(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right)+\epsilon_{2}\right)\wedge 1\right]$;
  3. (3)
    a remaining group of singular values in $[0,\epsilon_{2}\wedge 1]$.
Proof.
@@
$\sigma_{r}(\widehat{M})$ $=\sigma_{r}(P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}+P_{\operatorname*{col}(X_{1})}\Delta_{2}+\Delta_{1}P_{\operatorname*{col}(X_{2})}+\Delta_{1}\Delta_{2})$
@@
Note that by construction we have:
  1. (1)
    $(A_{1}+A_{2})^{\top}(A_{3}+A_{4})=0$,
  2. (2)
    $A_{1}A_{2}^{\top}=0$.
Then it holds that
$\sigma_{r}(\widehat{M})$ $\geq\sigma_{r}(A_{1}+A_{2})$ (Lemma A.1)
@@
Theorem A.3 (Estimation Error).
Suppose $\epsilon_{1}<\max\{1-\epsilon_{2},1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})-\epsilon_{2}\}$. Define ${R}_{\mathcal{J}}:=P_{\operatorname*{col}(X_{1})}\Delta_{2}+\Delta_{1}P_{\operatorname*{col}(X_{2})}+\Delta_{1}\Delta_{2}$ and $R_{\mathcal{I}_{k}}:=P_{\mathcal{J}}\Delta_{k}+\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}$, where $\Delta_{\mathcal{J}}:=P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}$ and $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$ Then, $\dim(\widehat{\mathcal{J}})=\dim({\mathcal{J}})$ and the deviation of the estimated joint subspace from the true joint subspace is bounded by
$\|P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}\|_{2}\leq\frac{\dim(\mathcal{J})^{1/2}\left\|R_{\mathcal{J}}+R^{\top}_{\mathcal{J}}\right\|_{2}}{1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})}.$
Furthermore, if the rank of each subspace view is correctly estimated, i.e. $\operatorname*{rank}(\widehat{X}_{k})=\operatorname*{rank}(X_{k})$, then, the deviation of the estimated individual subspace from the true individual subspace is bounded by
$\|P_{\widehat{\mathcal{I}}_{k}}-P_{\mathcal{I}_{k}}\|_{2}\leq 2\dim(\mathcal{I}_{k})^{1/2}\|R_{\mathcal{I}_{k}}\|_{F}.$
In the proof of Theorem 3.2, we use the adaptation of Davis-Kahan Theorem introduced in Yu et al. (2015). The theorem is stated in terms of the $\sin\Theta$ distance between the subspaces. To apply the theorem, we state the definition of the $\sin\Theta$ distance and principle angles here. Specifically, the $\sin\Theta$ distance is used to measure the difference between two $p\times r$ orthogonal columns $\widehat{V}$ and $V$. Suppose the singular values of $V^{\top}\widehat{V}$ are $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$. Then we call
$\Theta(\widehat{V},V)=\operatorname{diag}(\cos^{-1}(\sigma_{1}),\cos^{-1}(\sigma_{2}),\dots,\cos^{-1}(\sigma_{r}))$
as the principle angles. A quantitative measure of distance between the column spaces of $\widehat{V}$ and $V$ is then $\|\sin\Theta(\widehat{V},V)\|_{2}$ or $\|\sin\Theta(\widehat{V},V)\|_{F}$. Note that the difference of projection matrices, $(P_{\operatorname*{col}(\widehat{V}}-P_{\operatorname*{col}(V)})$, is related to the $\sin\Theta$ via the equality $\|P_{\operatorname*{col}(\widehat{V})}-P_{\operatorname*{col}({V}})\|_{2}$ = $\|\sin\Theta(\widehat{V},V)\|_{2}$
Proof.
@@
Joint components. To apply Davis-Kahan theorem, we show that there exits an eigengap in the population-level model $S=\tfrac{1}{2}\left(P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}+P_{\operatorname*{col}(X_{2})}P_{\operatorname*{col}(X_{1})}\right)$.
From Section 3.1, we know that $S=P_{\mathcal{J}}+\frac{1}{2}\left(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}+P_{\mathcal{N}_{2}}P_{\mathcal{N}_{1}}\right)$. Fix an orthonormal basis, $U_{\mathcal{J}}$, for $P_{\mathcal{J}}$. Since $\mathcal{J}\perp\mathcal{N}_{k}$, we have that $SU_{\mathcal{J}}=U_{\mathcal{J}}$. Hence, $S$ has at least $\dim(\mathcal{J})$ eigenvalues that are equal to $1$ with corresponding eigenvectors forming $\mathcal{J}$. Now fix any orthonormal vector $u\in\operatorname*{span}(U_{\mathcal{J}}^{\perp})$. Then we have:
@@
$\leq\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}$ (by triangle-inequality)
[…]
Therefore, the spectrum of $S$ becomes $\sigma_{r}(S)=1$ for $1\leq r\leq\dim(\mathcal{J})$ and $\sigma_{r}(S)\leq\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})<1$ for $r>\dim(\mathcal{J})$.
From here, applying Davis-Kahan, we have that:
$\|\sin\Theta(U_{\widehat{\mathcal{J}}},U_{\mathcal{J}})\|_{2}\leq\frac{2\dim(\mathcal{J})^{1/2}\|\widehat{S}-S\|_{2}}{1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})}$
@@
$\|P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}\|_{2}$ $=\|\sin\Theta(U_{\widehat{\mathcal{J}}},U_{\mathcal{J}})\|_{2}$
$\leq\frac{2\dim(\mathcal{J})^{1/2}\left\|\widehat{S}-S\right\|_{2}}{1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})}$

### v2

@@ 6.2. Nutrigenomic mice data
Figure 3. Left: Scree plot for gene expression and lipid content data. Blue point denotes the cutoff suggested by Gavish & Donoho (2014). Red point indicates our suggested cutoff. Middle / Right: Spectrum of the product of projection matrices onto each view for two different marginal ranks. The blue region indicates the random direction bound, $0,\lambda_{+}^{1/2}]$, used to filter out the noise directions. The green region is given by $1-\widehat{\varepsilon}_{1},1]$ where $\widehat{\varepsilon}_{1}$ is a bootstrap estimate of $\varepsilon_{1}$.
Table 3. Mean genotype and diet misclassification rates in percentages together with the minimum principle angle between estimated individual subspaces.
@@ Appendix A Techinical details
A.1. Proof of Theorem 3.1
Our proof relies on the following lemma, which we establish first.
@@
    $\sigma_{r}(X+Y)\geq\max(\sigma_{r}(X),\sigma_{r}(Y))$ for all $r\geq 1$ (assuming $\sigma_{1}\geq\sigma_{2}\geq....0$)
@@
$\lambda_{n}(X^{\top}X)+\lambda_{r}(Y^{\top}Y)$ $\leq\lambda_{r}(X^{\top}X+Y^{\top}Y)\leq\lambda_{1}(Y^{\top}Y)+\lambda_{r}(X^{\top}X)$
@@
$\lambda_{r}(X^{\top}X+Y^{\top}Y)\geq\max\left\{\lambda_{n}(X^{\top}X)+\lambda_{r}(Y^{\top}Y),\lambda_{n}(X^{\top}X)+\lambda_{r}(Y^{\top}Y)\right\}$
Recall that $\sigma^{2}_{r}(A)=\lambda_{r}(A^{\top}A)$. Since $\lambda_{n}(Y^{\top}Y)\geq 0$, $\lambda_{n}(X^{\top}X)\geq 0$, we get (1).
Similarly,
$\lambda_{r}(X^{\top}X+Y^{\top}Y)\leq\min\left\{\lambda_{r}(X^{\top}X)+\lambda_{1}(Y^{\top}Y),\lambda_{1}(X^{\top}X)+\lambda_{r}(Y^{\top}Y)\right\},$
since $\lambda_{1}(Y^{\top}Y)\geq 0$, $\lambda_{1}(X^{\top}X)\geq 0$, taking $r=1$ yields (2). ∎
of Theorem 3.1.
@@
[…]
@@
Note that by construction we have $(A_{1}+A_{2})^{\top}(A_{3}+A_{4})=0$ and $A_{1}A_{2}^{\top}=0$. Then it holds that:
[…]
@@
A.2. Proof of Theorem 3.2
In the proof of Theorem 3.2, we use the adaptation of the Davis-Kahan theorem as used in Yu et al. (2015) and Vershynin (2025). The theorem is stated in terms of the $\sin\Theta$ distance between the subspaces. To apply the theorem, we state the definition of the $\sin\Theta$ distance and principle angles here. Specifically, the $\sin\Theta$ distance is used to measure the difference between two $n\times r$ matrices $V$ and $\widehat{V}$ with orthogonal columns. Suppose the singular values of $V^{\top}\widehat{V}$ are $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$. Then we call
$\Theta(V,\widehat{V})=\operatorname{diag}(\cos^{-1}(\sigma_{1}),\cos^{-1}(\sigma_{2}),\dots,\cos^{-1}(\sigma_{r}))$
[…]
Lemma A.2 (Exercise 4.13 from Vershynin (2025)).
Let $\Sigma,\widehat{\Sigma}\in\mathbb{R}^{n\times n}$ be symmetric with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and $\widehat{\lambda}_{1}\geq\dots\geq\widehat{\lambda}_{n}$ respectively. Fix $1\leq s<n$ with $\lambda_{s}-\lambda_{s+1}>0$. Let $V=(v_{1},v_{2},\dots,v_{s})\in\mathbb{R}^{n\times s}$ and let $\widehat{V}=(\widehat{v}_{1},\widehat{v}_{2},\dots,\widehat{v}_{s})\in\mathbb{R}^{n\times s}$ consist of the first $s$ eigenvectors of $\Sigma$ and $\widehat{\Sigma}$. Then,
$\|\sin\Theta(V,\widehat{V})\|_{2}\leq\frac{2\|\Sigma-\widehat{\Sigma}\|_{2}}{\lambda_{s}-\lambda_{s+1}}.$
of Theorem 3.2.
@@
Joint components. To apply Lemma A.2, we show that there exits an eigengap in the population-level model $S=\tfrac{1}{2}\left(P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}+P_{\operatorname*{col}(X_{2})}P_{\operatorname*{col}(X_{1})}\right)$. The matrix $\widehat{S}=\tfrac{1}{2}\left(P_{{\operatorname*{col}}(\widehat{X}_{1})}P_{{\operatorname*{col}}(\widehat{X}_{2})}+P_{{\operatorname*{col}}(\widehat{X}_{2})}P_{{\operatorname*{col}}(\widehat{X}_{1})}\right)$ represents the finite sample estimate.
From Section 3.1, we know that $S=P_{\mathcal{J}}+\frac{1}{2}\left(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}+P_{\mathcal{N}_{2}}P_{\mathcal{N}_{1}}\right)$. Fix an orthonormal basis, $U_{\mathcal{J}}$, for $P_{\mathcal{J}}$. Since $\mathcal{J}\perp\mathcal{N}_{k}$, we have that $SU_{\mathcal{J}}=U_{\mathcal{J}}$. Hence, $S$ has at least $\dim(\mathcal{J})$ eigenvalues that are equal to $1$ with corresponding eigenvectors forming $\mathcal{J}$. Now fix any orthonormal vector $u\in\operatorname*{span}(U_{\mathcal{J}})^{\perp}$. Then we have:
@@
$\leq{\left\|P_{\mathcal{J}}u\right\|_{2}+\frac{1}{2}\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}(u)\right\|_{2}+\frac{1}{2}\left\|P_{\mathcal{N}_{2}}P_{\mathcal{N}_{1}}(u)\right\|_{2}}$ (by triangle-inequality)
$\leq\left\|P_{\mathcal{N}_{2}}P_{\mathcal{N}_{1}}\right\|_{2}$
$<1$ $\text{(since }\mathcal{I}_{1}\cap\mathcal{I}_{2}=\{\mathbf{0}\})$
Here, the second inequality follows from $u\in\mathcal{J}^{\perp}$, $\|u\|_{2}=1$ and the property that $\|Au\|_{2}\leq\|A\|_{2}\|u\|_{2}=\|A^{\top}\|_{2}\|u\|$ for any matrix $A\in\mathbb{R}^{n\times n}$. Therefore, the spectrum of $S$ becomes $\sigma_{i}(S)=1$ for $1\leq i\leq\dim(\mathcal{J})$ and $\sigma_{i}(S)\leq\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}<1$ for $i>\dim(\mathcal{J})$. Fix an orthonormal basis, $\widehat{U}_{\mathcal{J}}$, for $P_{\widehat{\mathcal{J}}}$. By the upper bound on $\varepsilon_{1}$ and appealing to Theorem 3.1, we have that $\mathrm{dim}(\widehat{\mathcal{J}})=\mathrm{dim}(\mathcal{J})$. Let $s:=\mathrm{dim}(\mathcal{J})$, $\widehat{\Sigma}=\widehat{S}$, and apply Lemma A.2 to obtain:
$\|\sin\Theta(U_{\mathcal{J}},U_{\widehat{\mathcal{J}}})\|_{2}\leq\frac{{2}\|S-\widehat{S}\|_{2}}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}$
@@
$\|P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}\|_{2}$ $=\|\sin\Theta(U_{\mathcal{J}},U_{\widehat{\mathcal{J}}})\|_{2}$
$\leq\frac{{2}\left\|S-\widehat{S}\right\|_{2}}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}$
$\leq\frac{{}\left\|P_{\operatorname*{col}(X_{1})}\Delta_{2}+\Delta_{1}P_{\operatorname*{col}(X_{2})}+\Delta_{1}\Delta_{2}+\Delta_{2}P_{\operatorname*{col}(X_{1})}+P_{\operatorname*{col}(X_{2})}\Delta_{1}+\Delta_{2}\Delta_{1}\right\|_{2}}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}$
$=\frac{{}\left\|R_{\mathcal{J}}+R_{\mathcal{J}}^{\top}\right\|_{2}}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}.$
Individual components. Since $\mathcal{J}\perp\mathcal{I}_{1}$ and $\mathcal{J}\perp\mathcal{I}_{2}$, we have that $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{1})}=P_{\mathcal{I}_{1}}$ and $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{2})}=P_{\mathcal{I}_{2}}$. Therefore, we have that $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{k})}$ has exactly $r_{\mathcal{I}_{k}}$ eigenvalues equal to $1$ and the rest equal to 0 so that the gap between the $r_{\mathcal{I}_{k}}$ and $r_{\mathcal{I}_{k}}+1$ eigenvalue of $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{k})}$ is one. Let $U_{{\mathcal{I}}_{k}}$ consist of the top $r_{\mathcal{I}_{k}}$ singular vectors of $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{1})}$: these span the subspace $\mathcal{I}_{k}$}. Let ${U}_{\widehat{\mathcal{I}}_{k}}$ consist of the first $\mathrm{rank}(\widehat{X}_{k})-\widehat{r}_{\mathcal{J}}$ singular vectors of $P_{\widehat{\mathcal{J}}^{\perp}}P_{\operatorname*{col}(\widehat{X}_{1})}$: these span the subspace $\widehat{\mathcal{I}}_{k}$. Since $\mathrm{rank}(\widehat{X}_{k})=\mathrm{rank}(X_{k})$ by assumption and $\widehat{r}_{\mathcal{J}}={r}_{\mathcal{J}}$ from the first part of the theorem, we have that $U_{{\mathcal{I}}_{k}}$ and $U_{\widehat{\mathcal{I}}_{k}}$ have the same number of columns. Let $s:=r_{\mathcal{I}_{k}}$, $\Sigma:=P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{k})}$, $\widehat{\Sigma}=P_{\widehat{\mathcal{J}}^{\perp}}P_{\operatorname*{col}(\widehat{X}_{1})}$, and apply Lemma A.2 to obtain:
$\|P_{\mathcal{I}_{k}}-P_{\widehat{\mathcal{I}}_{k}}\|_{2}$ $=\|\sin\Theta(U_{{\mathcal{I}_{k}}},U_{\widehat{\mathcal{I}}_{k}})\|_{2}$
$\leq{2\left\|P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{k})}-P_{\widehat{\mathcal{J}}^{\perp}}P_{\operatorname*{col}(\widehat{X}_{k})}\right\|_{2}}$
${=2\left\|P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{k})}-(P_{\mathcal{J}^{\perp}}+\Delta_{\mathcal{J}})(P_{\operatorname*{col}(X_{k})}-\Delta_{k})\right\|_{2}}$
${=2\left\|P_{{\mathcal{J}}^{\perp}}\Delta_{k}-\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}\right\|_{2}}$
${=2\|R_{\mathcal{I}_{k}}\|_{2}.}$
A.3. Proof of Theorem 3.3
The proof of this theorem relies on the following lemmas.
Lemma A.3.
Let $L\in\mathbb{R}^{d_{1}\times d_{2}}$ have entries that are independently and identically distributed as zero-mean Gaussian random variables with variance $\tau$. Then, with probability greater than $1-2\exp(-\max\{d_{1},d_{2}\})$, we have that $\|L\|_{2}\leq C\tau(d_{1}^{1/2}+d_{2}^{1/2})$, where $C$ is a constant (does not depend on $d_{1}$, $d_{2}$, or $\tau$).
Lemma A.3 follows from Theorem 4.4.3 of Vershynin (2025).
Lemma A.4.
Let $L\in\mathbb{R}^{d_{1}\times d_{2}}$ have entries that are independently and identically distributed as zero-mean Gaussian random variables with variance $\tau$. Let $U\in\mathbb{R}^{d_{1}\times t}$ and $V\in\mathbb{R}^{d_{2}\times t}$ be matrices with orthonormal columns with $t\leq\min\{d_{1},d_{2}\}$. Then $LV\in\mathbb{R}^{d_{1}\times t}$ and $L^{\top}{U}\in\mathbb{R}^{d_{2}\times t}$ are matrices with entries that are independent and identically distributed as zero-mean Gaussian random variables with variance $\tau$.
Lemma A.4 is a standard result based on the fact that linear combinations of Gaussian random variables are Gaussian. We state it as a separate lemma for convenience.
Lemma A.5.
Suppose Assumptions 12 hold. If $\widehat{r}_{k}=r_{k}$, then we have for some absolute constant $C>0$, with probability greater than $1-\mathcal{O}(\exp\{-\max(p_{k},n)\})$
$\|\Delta_{k}\|_{2}\leq C\left(\delta_{k}\frac{n^{1/2}+r_{k}^{1/2}}{\sigma_{r_{k}}(X_{k})}+\delta_{k}^{2}\frac{(n^{1/2}+p_{k}^{1/2})(p_{k}^{1/2}+r_{k}^{1/2})}{\sigma_{r_{k}}^{2}(X_{k})}\right).$
If $\widehat{r}_{k}\geq r_{k}$, then we have for some absolute constant $C>0$, with probability greater than $1-\mathcal{O}(\exp\{-\max(p_{k},n)\})$
$\|P_{\operatorname*{col}(X_{k})}\Delta_{k}\|_{2}\leq C\left(\delta_{k}\frac{n^{1/2}+r_{k}^{1/2}}{\sigma_{r_{k}}(X_{k})}+\delta_{k}^{2}\frac{(n^{1/2}+p_{k}^{1/2})(p_{k}^{1/2}+r_{k}^{1/2})}{\sigma_{r_{k}}^{2}(X_{k})}\right).$
of Lemma A.5.
We consider the two cases separately.
Correct rank estimation case with $\widehat{r}_{k}=r_{k}$ For simplicity, we drop subscript $k$ in the proof. Let $X$ of rank-$r$ have singular value decomposition
$X=[U\ U_{\perp}]\begin{pmatrix}\Sigma_{1}&0\\ 0&0\end{pmatrix}[V\ V_{\perp}]^{\top}.$
Let $Y=X+Z$, and similarly partition $Y$ as follows:
$Y=[\widehat{U}\ \widehat{U}_{\perp}]\begin{pmatrix}\widehat{\Sigma}_{1}&0\\ 0&\widehat{\Sigma}_{2}\end{pmatrix}[\widehat{V}\ \widehat{V}_{\perp}$$^{\top}$,
where $\widehat{U}\in\mathbb{R}^{n\times r},\widehat{V}\in\mathbb{R}^{p\times r}$, and $\widehat{X}=\widehat{U}\widehat{\Sigma}_{1}\widehat{V}^{\top}$. Then $\|\Delta\|_{2}=\|P_{{\operatorname*{col}(X)}}-P_{\operatorname*{col}(\widehat{X})}\|_{2}=\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2}=\|\widehat{U}_{\perp}^{\top}U\|_{2}$, where the last equality is true because of equality in ranks, e.g. Lemma 2.5 in Chen et al. (2021). Consider
$\widehat{U}_{\perp}^{\top}U$ $=\widehat{U}_{\perp}^{\top}U\Sigma_{1}V^{\top}V\Sigma_{1}^{-1}=\widehat{U}_{\perp}^{\top}(X)V\Sigma_{1}^{-1}=\widehat{U}_{\perp}^{\top}(Y-Z)V\Sigma_{1}^{-1}$
$=\widehat{U}_{\perp}^{\top}(\widehat{U}\widehat{\Sigma}_{1}\widehat{V}^{\top}+\widehat{U}_{\perp}\widehat{\Sigma}_{2}\widehat{V}_{\perp}^{\top}-Z)V\Sigma_{1}^{-1}$
$=\widehat{\Sigma}_{2}\widehat{V}_{\perp}^{\top}V\Sigma_{1}^{-1}-\widehat{U}_{\perp}^{\top}ZV\Sigma_{1}^{-1}.$
Similarly, we can show
$\widehat{V}_{\perp}^{\top}V=\widehat{\Sigma}_{2}\widehat{U}_{\perp}^{\top}U\Sigma_{1}^{-1}-\widehat{V}_{\perp}^{\top}Z^{\top}U\Sigma_{1}^{-1}.$
Combining the above two displays gives
$\widehat{U}_{\perp}^{\top}U$ $=\widehat{\Sigma}_{2}^{2}\widehat{U}_{\perp}^{\top}U\Sigma_{1}^{-2}-\widehat{\Sigma}_{2}\widehat{V}_{\perp}^{\top}Z^{\top}U\Sigma_{1}^{-2}-\widehat{U}_{\perp}^{\top}ZV\Sigma_{1}^{-1}.$
Therefore,
$\|\widehat{U}_{\perp}^{\top}U\|_{2}\leq\|\widehat{U}_{\perp}^{\top}U\|_{2}\|\widehat{\Sigma}_{2}\|_{2}^{2}\frac{1}{\sigma_{r}(X)^{2}}+\|\widehat{\Sigma}_{2}\|_{2}\frac{1}{\sigma_{r}(X)^{2}}\|\widehat{V}_{\perp}^{\top}Z^{\top}U\|_{2}+\frac{1}{\sigma_{r}(X)}\|\widehat{U}_{\perp}^{\top}ZV\|_{2}.$
From Weyl’s inequality and using that $X$ is rank-$r$,
$\|\widehat{\Sigma}_{2}\|_{2}\leq\|\Sigma_{2}\|_{2}+\|Z\|_{2}=\|Z\|_{2}.$
Therefore,
$\|\widehat{U}_{\perp}^{\top}U\|_{2}\leq\|\widehat{U}_{\perp}^{\top}U\|_{2}\frac{\|Z\|_{2}^{2}}{\sigma_{r}(X)^{2}}+\frac{\|Z\|_{2}}{\sigma_{r}(X)}\frac{\|\widehat{V}_{\perp}^{\top}Z^{\top}U\|_{2}}{\sigma_{r}(X)}+\frac{\|\widehat{U}_{\perp}^{\top}ZV\|_{2}}{\sigma_{r}(X)}$
If $\|Z\|_{2}\leq\frac{1}{2^{1/2}}\sigma_{r}(X)$, then
$\|\widehat{U}_{\perp}^{\top}U\|_{2}\leq 2\left\{\frac{\|\widehat{U}_{\perp}^{\top}ZV\|_{2}}{\sigma_{r}(X)}+\frac{\|Z\|_{2}}{\sigma_{r}(X)}\frac{\|\widehat{V}_{\perp}^{\top}Z^{\top}U\|_{2}}{\sigma_{r}(X)}\right\}.$
Appealing to Lemma A.3, we have that $\|Z\|_{2}\leq C_{1}\delta(n^{1/2}+p^{1/2})$ with probability greater than $1-2\exp(-\max\{n,p\})$ for some constant $C_{1}$. Thus, under the assumption that $\sigma_{r}(X)\geq\sqrt{2^{1/2}}C_{1}\delta(n^{1/2}+p^{1/2})$, we get that $\|Z\|_{2}\leq\frac{1}{2^{1/2}}\sigma_{r}(X)$ holds with high probability. Furthermore, we have with probability greater than $1-2\exp(-\max\{n,p\})$
$\begin{aligned} \|\widehat{U}_{\perp}^{\top}ZV\|_{2}&\leq\|ZV\|_{2}\leq C_{2}\delta(n^{1/2}+r^{1/2}),\\ \|\widehat{V}_{\perp}^{\top}Z^{\top}{U}\|_{2}&\leq\|Z^{\top}{U}\|_{2}\leq C_{3}\delta(p^{1/2}+r^{1/2}),\end{aligned}$
where $C_{2}$ and $C_{3}$ are constants that do not depend on $n,p,r$ and $\delta$. The first inequality in each of the equations above is due to the sub-multiplicative property of the spectral norm, and the second inequality follows by combining Lemma A.4 with Lemma A.3. Putting things together and bringing back the notation for every view, we have that with probability greater than $1-\mathcal{O}(-\max\{n,p_{k}\})$,